# Heterogeneity in Entity Matching: A Survey and Experimental Analysis

### v1

@@ 3.7 LLMs
LLMs are increasingly prominent in entity matching (EM), offering strong capabilities for handling heterogeneity. Their ability to generalize across tasks and domains makes them highly adaptable to diverse data types and formats. With zero-shot and few-shot learning, LLMs can operate effectively in low-supervision settings—a critical advantage in HEM scenarios where labeled data is often limited. LLMs leverage pre-trained knowledge to align records across differing schemas and semantic contexts, enabling robust performance on complex matching tasks.
@@
Beyond text, LLMs can reason over multimodal data by integrating heterogeneous input types such as structured tables, numerical fields, and free-text. Retrieval-augmented methods allow dynamic access to external knowledge sources, helping the model adapt to domain-specific vocabulary or evolving context. LLMs can also synthesize training data or provide natural language explanations for match decisions, enhancing both performance and interpretability. These characteristics make them particularly suited for the diverse challenges of heterogeneity in EM.
Recent work explores a range of LLM-based EM strategies. MatchGPT [59] examines ChatGPT (gpt3.5-turbo-0301) in zero-shot and in-context settings. It finds that ChatGPT can perform competitively with minimal supervision, particularly when guided by well-designed prompts and matching rules. BoostER [37] improves the cost-efficiency of LLM-driven EM by selecting questions that minimize uncertainty, using entropy-based heuristics to balance label quality and API cost.
FT-LLM [71] investigates the fine-tuning of LLMs for EM, showing that structured explanations and example selection significantly enhance in-domain performance, especially for smaller models. However, generalization to new domains remains a challenge. COMEM [83] proposes a three-pronged framework that leverages LLMs for matching, comparing, and selecting. The “selecting” strategy, which integrates global record context, is particularly effective in complex HEM tasks that require consistency across large datasets.
@@
Complementing these efforts, [54] analyze the internal behavior of BERT-based matchers, studying attention stability, token influence, and sensitivity to input order. Their goal is interpretability rather than performance. [55] extend this work by evaluating additional datasets and perturbations, revealing how domain shift, sequence length, and training data size affect transformer-based EM.
@@
Collectively, these studies highlight the versatility of LLMs in EM, and their promise for addressing the multifaceted challenges of HEM—particularly through prompt engineering, adaptation, and integration with external knowledge sources.
@@ 4 Entity Matching and the FAIR Principles
The FAIR principles promote data practices that make information Findable, Accessible, Interoperable, and Reusable [85]. These guidelines are now widely embraced across scientific and industrial domains to support data sharing, reproducibility, and integration. Central to this vision is the need to link and reuse datasets across organizational and technological boundaries—a task often complicated by the pervasive heterogeneity of real-world data.
As we have shown, HEM refers to the challenges EM systems face when dealing with representation and semantic differences across datasets. These same heterogeneities pose barriers to FAIRification. Conversely, robust EM techniques—particularly those designed to handle heterogeneity—are essential to realizing FAIR goals. In this section, we unpack this dual relationship: how FAIR principles guide EM under heterogeneity, and how HEM complicates the path to FAIR compliance.
  • Findability. Findability requires that datasets and metadata be reliably indexed and retrievable by humans and machines, using persistent identifiers and rich, structured metadata. In heterogeneous settings, EM contributes by resolving duplicated or fragmented records and standardizing identifiers across formats. For example, provenance-aware EM pipelines help maintain traceability across sources, a key requirement for persistent findability. However, representation heterogeneity—such as nested JSON in one source versus flat tables in another—can obscure metadata extraction and record indexing. Semantic heterogeneity further complicates discovery when entities are described using inconsistent or ambiguous terms, reducing the effectiveness of search-based retrieval and downstream matching.
  • Accessibility. Accessibility ensures that data and metadata are retrievable under well-defined protocols, with clear licensing and machine-readable formats. From an EM perspective, accessibility is improved when schemas are consistently documented and attributes are harmonized—tasks that often depend on schema matching and terminology alignment. Methods such as GraphER [36] and DITTO [40], which handle schema-free or loosely structured inputs, enable flexible access across diverse data representations. Yet, these same structural mismatches hinder accessibility by creating barriers to uniform interpretation, especially when schemas are undocumented or inconsistently applied. Semantic variation adds a second layer of difficulty: identical concepts labeled differently across sources challenge even well-designed APIs and access protocols.
  • Interoperability. Interoperability aims to ensure that systems and datasets can be combined and understood in a unified manner, relying on shared vocabularies, ontologies, and data models. HEM directly obstructs this goal. For instance, schema heterogeneity requires EM systems to align fields with differing structures or granularity. Ontology-aware methods, such as OGKs [43], explicitly target this challenge by using external knowledge to reconcile semantic mismatches. However, full interoperability remains difficult when context-dependent meanings or evolving domain terms cannot be resolved via static vocabularies. EM methods must incorporate contextual embeddings and hierarchical alignment—e.g., as in HierGAT [87]—to approximate interoperability in the presence of complex heterogeneity.
  • Reusability. Reusability focuses on enabling future applications of data through clear documentation, quality assurance, and contextual information. High-quality entity matching improves reusability by resolving duplicate records, enriching metadata, and standardizing semantics. For example, contrastive learning-based approaches such as Sudowoodo [82] enhance latent semantic clustering, making entities easier to reuse across datasets. Nevertheless, HEM undermines reusability when records are noisy, incomplete, or inconsistently annotated. Poor attribute-level matching or unresolved ambiguities in entity identity can propagate downstream, reducing the trustworthiness and utility of integrated datasets. Improving reusability thus requires EM methods to be robust not only to heterogeneity but also to noise and uncertainty.
FAIR principles provide a valuable lens for diagnosing the challenges of heterogeneity in entity matching. At the same time, HEM-aware EM systems—especially those that use schema-agnostic architectures, semantic embeddings, and provenance tracking—can actively contribute to building FAIR-compliant data infrastructures. Closing this loop between EM and FAIRification is essential for creating transparent, reusable, and trustworthy data systems.
@@ 5.1.1 Datasets and Preparation
We use six widely studied datasets—Abt-Buy, Company, Fodor-Zagat, WDC, Walmart-Amz, and iTunes-Amz —summarized in Table 1. These datasets span diverse characteristics: small to large sizes (from hundreds to hundreds of thousands of records), structured and textual attributes, and hierarchical fields (e.g., “category” and “brand” in Walmart-Amz). This diversity allows us to apply all experimental perturbations described in the following sections. Including additional datasets or both clean and noisy variants would expand the study without altering the core findings.
@@
  • Synonym Injection: To simulate semantic heterogeneity from terminology variation, we replace words in textual attributes with contextually appropriate synonyms. This is applied to datasets with rich textual content—Abt-Buy, Company, and WDC. We first extract words from the textual attributes, removing stopwords, numeric tokens, and non-alphabetic terms. We use the KeyBERT library [29] to extract candidate keywords, filtering out product names and domain-specific terms.
@@
    To generate synonyms, we compare WordNet [45], BERT [12], and LLMs such as GPT-4 and Gemini against a small manually labeled test set. GPT-4 outperforms all others in contextual accuracy, so we adopt it to generate synonyms. Using a prompt-based approach, we ask GPT-4 to replace words with their most appropriate synonyms based on sentence context. These replacements are applied randomly to a specified proportion of candidate words. At 100% synonym ratio, we modify approximately 440k/3.9M tokens in WDC, 257k/616k in Abt-Buy, and 84M/237M in Company (test/train).
  • Hierarchical Data Distortion: To simulate granularity-based semantic heterogeneity, we modify hierarchical attributes such as time and location. For location, we alter granularity by replacing “city” values with “state” or “country” (or vice versa). For time, we convert between formats such as “hours” and “days.” These changes introduce controlled distortions in a subset of attributes, varying the level of aggregation.
@@
    We quantify information loss using entropy, computed as the sum of individual column entropies across independent attributes, following standard information theory [66]. Datasets with suitable hierarchies—iTunes-Amz, Walmart-Amz, and Fodor-Zagat —are selected for these experiments. The number of affected test/train cells is approximately 450/1.7k for iTunes-Amz, 550/2.2k for Fodor-Zagat, and 6.1k/24.5k for Walmart-Amz.
  • Dirty Data Injection: To simulate data quality heterogeneity, we introduce missing values, attribute noise, and label noise. For missing values, we follow three standard patterns. In the Missing Completely at Random (MCAR) setting, values are removed uniformly without any dependency on the data. In the Missing at Random (MAR) setting, missingness is conditioned on observed attributes—for example, price values might be removed only for products in specific categories. In the Missing Not at Random (MNAR) setting, missingness depends on the unobserved value itself, such as selectively removing high prices.
    For attribute noise, values are replaced with unrelated random values. Label noise is introduced by flipping match labels for a subset of record pairs based on a fixed noise ratio. Further injection details, including which attributes are affected under each condition, are documented in our repository [47].

### v2

@@ 3.13 Large Language Models
LLMs have recently emerged as powerful tools for EM [94, 66, 113, 133, 88, 89], offering strong capabilities for handling heterogeneity. Their ability to generalize across tasks and domains makes them highly adaptable to diverse data types and formats. With zero-shot and few-shot learning, LLMs can operate effectively in low-supervision settings—a critical advantage in HEM scenarios where labeled data is often limited. LLMs leverage pre-trained knowledge to align records across differing schemas and semantic contexts, enabling robust performance on complex matching tasks.
@@
Beyond text, LLMs can reason over multimodal data by integrating heterogeneous input types such as structured tables, numerical fields, and free text. Retrieval-augmented methods allow dynamic access to external knowledge sources, helping the model adapt to domain-specific vocabulary or evolving context. LLMs can also synthesize training data or provide natural-language explanations for match decisions, enhancing both performance and interpretability. These characteristics make them particularly suited for the diverse challenges of heterogeneity in EM.
@@
Early work such as [88] analyzes the internal behavior of BERT-based matchers, studying attention stability, token influence, and sensitivity to input order, with a primary focus on interpretability. An extended study by [89] evaluates additional datasets and perturbations, revealing how domain shift, sequence length, and training data size affect transformer-based EM.
More recent work directly applies LLMs to EM tasks. MatchGPT [94] evaluates ChatGPT (gpt3.5-turbo-0301) in zero-shot and in-context settings, showing that competitive performance is possible with carefully designed prompts and rules. BoostER [66] improves the cost-efficiency of LLM-driven EM by selecting questions that minimize uncertainty using entropy-based heuristics, balancing annotation quality and API cost.
FT-LLM [113] investigates fine-tuning strategies for LLMs in EM, showing that structured explanations and example selection significantly boost in-domain accuracy, especially for smaller models. However, challenges remain in cross-domain generalization. COMEM [133] proposes a three-pronged framework—matching, comparing, and selecting—that leverages global record context. The “selecting” module, which integrates broader dataset-level semantics, is particularly effective in complex HEM scenarios.
@@
Collectively, these studies highlight the versatility of LLMs in EM and their promise for addressing the multifaceted challenges of HEM—particularly through prompt engineering, adaptation, interpretability, and integration with external knowledge sources.
3.14 Relating Surveyed Approaches to the HEM Taxonomy
Figure 1 organizes heterogeneity into eight second-level categories spanning representation and semantic differences. To make this taxonomy more operational, we briefly summarize how each topic reviewed in this section targets specific heterogeneity types. This mapping also clarifies more about why these methodological areas are relevant to HEM.
  • Schema and Structural Heterogeneity. Methods in Section 3.1 (e.g., HERA, Machamp, GraphER) primarily address schema/structural heterogeneity: mismatched attribute names, missing fields, varying nesting layouts, and inconsistent relational structure. They also interact with datatype/format heterogeneity (e.g., JSON vs. tables) and, to a lesser extent, granularity heterogeneity when attributes differ in resolution or abstraction across datasets.
  • Representation Learning and Semantic Embeddings. Section 3.2 targets terminology/vocabulary heterogeneity through contextual embeddings (e.g., DITTO, HierGAT). These models also mitigate contextual semantics by encoding relational and positional cues; handle within-dataset representation heterogeneity through noise-tolerant embedding spaces; and partially support datatype/format heterogeneity by operating over linearized or text-derived representations.
  • Knowledge Graphs and Ontologies for EM. Section 3.4 leverages ontologies and KGs to resolve terminology and contextual heterogeneity via hierarchical types, semantic relations, and synonym mappings. KGs also reduce schema/structural and granularity heterogeneity by offering schema-independent, logically grounded representations; and support limited linguistic heterogeneity through multilingual knowledge bases.
  • Ontology Matching and KG Alignment. Section 3.5 primarily addresses schema/structural and terminology heterogeneity by aligning classes, relations, and identifiers across heterogeneous ontologies. Modern KG alignment methods additionally target contextual semantics through structure-aware embeddings and handle linguistic heterogeneity through cross-lingual mappings and multilingual encoders.
  • Benchmarks for HEM. Section 3.6 presents datasets that instantiate multiple forms of heterogeneity simultaneously. These include representation heterogeneity (schema drift, datatype inconsistencies, granularity mismatches, multimodal attributes), semantic heterogeneity (terminology variation, contextual ambiguity), and linguistic heterogeneity (multilingual product descriptions and KG labels). Such benchmarks allow controlled evaluation across several dimensions of HEM.
  • Instance Coreference Resolution. Section 3.7 corresponds mainly to terminology, contextual, and linguistic heterogeneity: CR systems must reconcile aliases, abbreviations, and context-dependent references across documents and domains. CR also touches on granularity heterogeneity (e.g., entity vs. sub-entity mentions) and mild within-dataset representation heterogeneity when mentions vary in completeness or surface form.
  • Transfer Learning and Domain Adaptation. Section 3.8 addresses terminology, contextual, and linguistic heterogeneity by adapting models across domains with differing vocabularies, styles, and label distributions. Domain adaptation also mitigates datatype/format shifts and within-dataset representation heterogeneity by aligning feature distributions or learning domain-invariant representations.
  • Active Learning and Interactive Methods. Section 3.9 primarily helps resolve terminology and contextual heterogeneity by allowing models to query users on ambiguous or domain-specific pairs. Interactive correction also reduces the effects of granularity heterogeneity (e.g., attribute grouping differences) and within-dataset representation heterogeneity (inconsistent fields, missing values).
  • Progressive and Incremental EM. Section 3.10 relates primarily to within-dataset representation and schema/structural heterogeneity. Progressive ER methods reorder or prioritize comparisons when attribute overlap is low, schemas differ across sources, or similarity signals are unreliable. They also address datatype/format heterogeneity by adapting to partially missing or inconsistently typed attributes, and indirectly handle contextual heterogeneity by allocating computation to pairs with ambiguous or conflicting evidence.
  • Self-Supervised and Pseudo-Label EM. Section 3.11 directly targets terminology/vocabulary and contextual heterogeneity by learning representations from raw text, relational structure, or token–attribute–tuple graphs without needing aligned schemas or labeled examples. These approaches also mitigate within-dataset representation heterogeneity (noise, missing values, inconsistent fields) and support datatype/format heterogeneity by extracting domain-agnostic structure from mixed or semi-structured inputs. Contrastive and graph-based self-supervision additionally helps address mild schema/structural heterogeneity when attribute layouts differ across sources.
  • Evolutionary, Meta-Heuristic, and Hybrid Approaches. Section 3.12 addresses schema/structural, granularity, and within-dataset representation heterogeneity by evolving matching rules, feature subsets, or hybrid similarity operators. These systems also handle datatype/format heterogeneity through flexible rule search and occasionally incorporate contextual or terminology cues when combined with neural or symbolic components.
  • LLM-Based EM. Section 3.13 naturally handles terminology, contextual, and linguistic heterogeneity through pretrained semantic knowledge and in-context reasoning. LLMs also partially address datatype/format heterogeneity by interpreting semi-structured inputs (e.g., JSON, tables) as text, and can mitigate within-dataset representation heterogeneity through robust contextualization of noisy attributes.
This mapping shows that each methodological area in Section 3 addresses a distinct subset of heterogeneity challenges and collectively spans all eight second-level dimensions in our taxonomy. It also clarifies how the taxonomy guides the organization and interpretation of the surveyed literature.
@@ 4 Entity Matching and the FAIR Principles
The FAIR principles promote data practices that make information Findable, Accessible, Interoperable, and Reusable [136]. These guidelines are widely embraced across scientific and industrial domains to support data sharing, reproducibility, and large-scale data integration. Central to this vision is the ability to identify, link, and reuse entities across datasets that differ in structure, representation, and semantics—precisely the setting addressed by HEM.
As discussed throughout Section 3, representation and semantic heterogeneity simultaneously obstruct FAIRification and motivate the design of robust EM techniques. Schema mismatches, inconsistent representations, terminology variation, and granularity differences all complicate the realization of FAIR goals. Conversely, EM systems that are explicitly designed to operate under heterogeneity play a critical role in enabling FAIR-compliant data infrastructures.
To provide a more structured synthesis of this relationship, we explicitly connect the EM research areas reviewed in Section 3 to the individual FAIR principles. Table 1 summarizes how representative methods from each surveyed area contribute to Findability, Accessibility, Interoperability, and Reusability. Rows correspond directly to the major EM areas discussed in Section 3, and each cell lists concrete methods (with references) that most directly support a given FAIR dimension. The table is intended to highlight traceable technical connections rather than provide exhaustive coverage.
EM Area Findable Accessible Interoperable Reusable
Schema & structural heterogeneity (§3.1) HERA [70] GraphER [65] Machamp [129]
Repr. learning & sem emb (§3.2) Ditto [69] Ditto [69] HierGAT [141] Sudowoodo [132] Unicorn [124]
DL & GNNs (§3.3) Seq2SeqMatcher [85] LinKG [146] HierMatcher [37] R-SupCon [93] GTA [29]
KGs & ontologies for EM (§3.4) OGKs [72] Bornemann et al. [11]
Ontology matching & KG alignment (§3.5) LogMap [53] AML [52] PARIS [114] MTransE [15] BootEA [118] RDGCN [138] CrossKG [122] RREA [74] MultiKE [47] CollectiveEA [145]
Self-supervised & pseudo-label EM (§3.11) CollaborER [41] EmbDI [14] Sudowoodo [132]
Evolutionary & hybrid approaches (§3.12) ERGP [116] DeepER/DeeperFlow [42] DeepBlocker [119]
LLMs (§3.13) PromptEM [131] MatchGPT [94] COMEM [133]
Table 1: Structured mapping between EM research areas reviewed in Section 3 and the FAIR principles they most directly support. Each cell lists representative methods and references, highlighting concrete technical mechanisms rather than exhaustive coverage.
  • Findability. Findability requires that entities and metadata be consistently indexed and retrievable by humans and machines using persistent identifiers and well-defined representations. In heterogeneous settings, duplicated or fragmented entity descriptions hinder reliable indexing and discovery. EM approaches addressing schema and representation heterogeneity (Sections 3.1 and 3.2) directly support Findability by constructing normalized or schema-independent entity representations. For example, HERA [70] incrementally builds super-records across heterogeneous schemas, while contextual embedding models such as Ditto [69] produce canonical textual representations that enable consistent indexing across sources, as summarized in Table 1.
  • Accessibility. Accessibility emphasizes that data and metadata should be retrievable through well-defined, machine-readable protocols. Structural and format heterogeneity directly obstruct accessibility when schemas are undocumented, inconsistent, or incompatible across sources. Schema-agnostic EM methods reviewed in Sections 3.1 and 3.3—such as GraphER [65] and Ditto [69]—enable record comparison without requiring strict prior schema alignment. By operating over loosely structured or heterogeneous inputs, these approaches operationalize FAIR Accessibility in settings where explicit schema harmonization is infeasible (Table 1).
  • Interoperability. Interoperability aims to ensure that datasets can be combined and interpreted within a shared semantic framework, typically relying on common vocabularies, ontologies, or data models. Semantic heterogeneity—arising from terminology variation, granularity mismatches, and contextual ambiguity—poses a primary barrier to this goal. Ontology- and knowledge-aware EM methods reviewed in Sections 3.4 and 3.5, such as OGKs [72], PARIS [114], and RDGCN [138], explicitly encode semantic relations and hierarchical structure to reconcile mismatched meanings across datasets. Deep hierarchical models such as HierGAT [141] further integrate semantic and structural signals, supporting interoperability at both the schema and instance levels, as reflected in Table 1.
  • Reusability. Reusability focuses on enabling future use of data through semantic clarity, quality assurance, and trustworthiness. Unresolved heterogeneity in entity identity propagates ambiguity to downstream applications, reducing confidence in reuse. Self-supervised and pseudo-label EM methods (Section 3.11), such as Sudowoodo [132] and EmbDI [14], learn robust representations that generalize across heterogeneous datasets with limited supervision. Evolutionary and hybrid EM approaches (Section 3.12) further enhance reusability by combining neural, symbolic, and rule-based signals to produce more interpretable and adaptable matching outcomes (Table 1).
Viewed through the FAIR lens, heterogeneity is not merely an obstacle for entity matching but a defining constraint that shapes data reuse at scale. Conversely, the EM techniques surveyed in Section 3—particularly schema-agnostic models, semantic embeddings, knowledge-aware approaches, self-supervised methods, and hybrid pipelines—constitute concrete technical enablers of FAIRification. By making these connections explicit, Table 1 clarifies how advances in heterogeneity-aware EM directly support the construction of findable, accessible, interoperable, and reusable data infrastructures.
@@ 5.1.1 Datasets and Preparation
We use six widely studied datasets–Abt-Buy, Fodor-Zagat, WDC, Walmart-Amz, and iTunes-Amz–summarized in Table 2. These datasets span diverse characteristics: small to large sizes (from hundreds to hundreds of thousands of records), structured and textual attributes, and hierarchical fields (e.g., “category” and “brand” in Walmart-Amz). This diversity allows us to apply all experimental perturbations described in the following sections. Including additional datasets or both clean and noisy variants would expand the study without altering the core findings.
@@
  • Synonym Injection: To simulate semantic heterogeneity from terminology variation, we replace words in textual attributes with contextually appropriate synonyms. This is applied to datasets with rich textual content–Abt-Buy, and WDC. We first extract words from the textual attributes, removing stopwords, numeric tokens, and non-alphabetic terms. We use the KeyBERT library [45] to extract candidate keywords, filtering out product names and domain-specific terms.
@@
    To generate synonyms, we compare WordNet [78], BERT [22], and LLMs such as GPT-4 and Gemini against a small manually labeled test set. GPT-4 outperforms all others in contextual accuracy, so we adopt it to generate synonyms. Using a prompt-based approach, we ask GPT-4 to replace words with their most appropriate synonyms based on sentence context. These replacements are applied randomly to a specified proportion of candidate words. At 100% synonym ratio, we modify approximately 440k/3.9M tokens in WDC, 257k/616k in Abt-Buy, and 84M/237M in test/train).
    To rigorously distinguish between the effects of semantic heterogeneity and random lexical noise, we establish a Random Word Noise baseline. As implemented in our benchmarking scripts, this baseline isolates the identical token positions targeted by the synonym injection process. However, instead of using contextually relevant synonyms, we replace these tokens with unrelated words randomly sampled from the NLTK English word corpus. This process includes preprocessing steps to preserve structural consistency, such as handling compound terms (e.g., standardizing “light-emitting diode”) before injection. By maintaining the exact same noise distribution and token indices as the synonym set, this baseline allows us to attribute performance drops specifically to semantic drift rather than simple vocabulary mismatch.
  • Hierarchical Data Distortion: To simulate granularity-based semantic heterogeneity, we modify hierarchical attributes such as time, location, and categorization. Instead of random noise, we employ domain-specific taxonomy trees to systematically alter the level of abstraction. We implement two specific perturbation mechanisms based on the attribute type:
    • Categorical Generalization: For nominal attributes (e.g., City, Brand, Category), we utilize nested dictionaries to map specific entities to their semantic parents. For example, in the Walmart-Amz dataset, a specific brand like “Acer” is mapped to “Computers,” which is further mapped to “Electronics.” The perturbation function traverses this hierarchy to replace a leaf node with a randomly selected ancestor.
    • Numerical and Temporal Discretization: For continuous or high-cardinality values, we apply interval-based binning hierarchies. Exact values are replaced with range descriptors or broader timeframes. For instance, in the iTunes-Amz dataset, a specific song duration (e.g., 210 seconds) is generalized to a “Moderate” length bucket, while release dates are abstracted to their release year or decade (e.g., “1999” becomes “1990s”).
@@
    These transformations allow us to inject controlled semantic heterogeneity, simulating data integration scenarios where sources report at different levels of granularity. We quantify information loss using entropy, computed as the sum of individual column entropies across independent attributes, following standard information theory [106]. Datasets with suitable hierarchies–iTunes-Amz, Walmart-Amz, and Fodor-Zagat–are selected for these experiments. The number of affected test/train cells is approximately 450/1.7k for iTunes-Amz, 550/2.2k for Fodor-Zagat, and 6.1k/24.5k for Walmart-Amz.
  • Dirty Data Injection: To simulate data quality heterogeneity, we introduce missing values, attribute noise, and label noise. We implement distinct perturbation logic for each category to model real-world data corruption patterns:
    • Missing Values: We employ three mechanisms defined by their dependency structures.
        For Missing Completely at Random (MCAR), we uniformly sample row and column indices across the dataset to remove values, ensuring no dependency on the data content.
        For Missing at Random (MAR), the missingness probability is conditioned on the record’s ground truth label. We assign a higher base probability weight to matching record pairs (0.8) compared to non-matches (0.2). This weight is then passed through an arctangent function to generate a smoothed probability ($P=\arctan(weight)/\arctan(N)$), determining whether a value is masked.
        For Missing Not at Random (MNAR), we introduce a dependency on the unobserved value itself. The probability weight is calculated by combining the class label weight with a normalized hash of the attribute’s specific value (added as a factor $hash(value)\%100/100$). This ensures that specific values (e.g., specific price points or high-cardinality strings) have distinct probabilities of being missing.
    • Attribute Noise: We apply type-specific distortions.
        For string attributes, we simulate typographical errors using a two-step process: first, we select approximately 30% of the string’s positions and replace the characters with random ASCII letters; second, we append an additional random character to the end of the string to simulate insertion errors.
        For numerical attributes, we apply multiplicative noise by perturbing the original value with a random factor drawn uniformly from the range $20\%,+20\%]$ and rounding the result to the nearest integer.
    • Label Noise: To simulate annotation errors, we randomly select a subset of training and validation records determined by the noise ratio (e.g., 5–25%) and flip their binary labels ($0\leftrightarrow 1$), creating valid-but-incorrect supervision signals.